# Beyond Rate Coding: Surrogate Gradients Enable Spike Timing Learning in Spiking Neural Networks

### v2

@@ Abstract
  • The surrogate gradient descent algorithm enabled spiking neural networks to be trained to carry out challenging sensory processing tasks, an important step in understanding how spikes contribute to neural computations. However, it is unclear the extent to which these algorithms fully explore the space of possible spiking solutions to problems. We investigated whether spiking networks trained with surrogate gradient descent can learn to make use of information that is only encoded in the timing and not the rate of spikes. We constructed synthetic datasets with a range of types of spike timing information (interspike intervals, spatio-temporal spike patterns or polychrony, and coincidence codes). We find that surrogate gradient descent training can extract all of these types of information. In more realistic speech-based datasets, both timing and rate information is present. We therefore constructed variants of these datasets in which all rate information is removed, and find that surrogate gradient descent can still perform well. We tested all networks both with and without trainable axonal delays. We find that delays can give a significant increase in performance, particularly for more challenging tasks. To determine what types of spike timing information are being used by the networks trained on the speech-based tasks, we test these networks on time-reversed spikes which perturb spatio-temporal spike patterns but leave interspike intervals and coincidence information unchanged. We find that when axonal delays are not used, networks perform well under time reversal, whereas networks trained with delays perform poorly. This suggests that spiking neural networks with delays are better able to exploit temporal structure. Since humans also perform poorly with time-reversed speech, it suggests that these networks may be better suited as neural models. To facilitate further studies of temporal coding, we have released our modified speech-based datasets.
@@ 1 Introduction
Spiking Neural Networks (SNNs) are biologically inspired models that compute via discrete, sparse spikes, rather than continuous activations (Maass, 1997). This event‑driven framework not only captures rich temporal patterns (such as inter‑spike intervals and cross‑neuron synchrony) but also powers energy‑efficient neuromorphic hardware (Painkras et al., 2013; Akopyan et al., 2015; Davies et al., 2018; Orchard et al., 2021; Su et al., 2024).
@@
Despite their promise, a persistent challenge remains: how to convert precise spike timing into effective learning signals. Surrogate Gradient Descent (Surrogate GD) provides a practical solution by replacing the non‑differentiable spike with a smooth surrogate, thereby allowing backpropagation through time (Neftci et al., 2019). Subsequent work has improved its efficiency and adaptiveness by enforcing sparser gradients (Perez-Nieves and Goodman, 2021) and by developing more adaptive surrogate functions (Wang et al., 2023). Surrogate GD–trained networks have matched or outperformed rate‑based recurrent neural networks on several neuromorphic benchmarks (Shrestha and Orchard, 2018; Sun et al., 2025a; Dampfhoffer and Mesquida, 2024; Hammouamri et al., 2024; Sun et al., 2025b), but it remains unclear whether these successes stem from genuine temporal coding or from reliance on coarse firing‑rate features.
@@
To address this ambiguity, researchers have injected various temporal inductive biases: heterogeneous time constants (Fang et al., 2021; Perez-Nieves et al., 2021; Habashy et al., 2024); learnable axonal delays (Shrestha et al., 2022; Sun et al., 2022; Patiño-Saucedo et al., 2023); synaptic delays (Zhang et al., 2020; Yu et al., 2022; Sun et al., 2024; Hammouamri et al., 2024); and dendritic delay models (D’agostino et al., 2024; Zheng et al., 2024). Yet even with these enhancements, there has been no systematic test of whether or not Surrogate GD finds networks that make use of precise timing.
@@
This uncertainty leads to a central question: can Surrogate GD-trained SNNs learn and depend on spike timing codes beyond rate statistics? This question is important because if they were just using rate-based information, then (a) they are not making use of all the information that is available, and are therefore less efficient than they could be, (b) it limits their effectiveness as a model of spike-based computation in the brain, and (c) it raises the question of whether or not there is an advantage of using such expensive training methods compared to simply approximating rate-based networks. To address this central question, we used both synthetic datasets as well as more complicated real-world benchmarks. First, we construct synthetic benchmark tasks that isolate the contribution of inter-spike interval (ISI), coincidence, and cross-channel inter-spike interval (CCISI, a type of spatio-temporal spike pattern or polychrony) under controlled spike counts, building upon prior work that highlights the ability of SNNs to leverage temporal coding for classification (Bohte et al., 2002; Mirsadeghi et al., 2021). Then, we evaluate models on datasets with real-world temporal complexity (Cramer et al., 2020), introducing spike count-normalized variants to remove rate confounds. We assess robustness under biologically inspired perturbations: Gaussian jitter applied per-spike and per-neuron (Amarasingham et al., 2012; Agmon, 2012), and spike deletions (Mainen and Sejnowski, 1995; Chen et al., 2022). Finally, we analyse the effect of time reversal that selectively disrupts spike patterns while leaving coincidences and ISIs unchanged.
@@
Our findings show that Surrogate GD-trained SNNs can extract fine-grained temporal structure even without spike rate information, including intra-neuron ISIs, cross-neuron synchrony or coincidence, and cross-channel causal ISI. Delay learning further enhances the model’s ability to exploit such timing features, especially for long-range causal gaps that exceed the membrane integration window controlled by the time constant. Across perturbations we observe distinct degradation patterns that reveal both vulnerabilities and robustness in the model’s temporal encoding. Notably, under time reversal, Surrogate GD-trained SNNs not only leverage reversal-invariant features like ISI and coincidence, but also capture causal dependencies in spike timing (as in CCISI), indicating a layered temporal representation.
@@ 2.1 Models
Training is implemented in the SLAYER framework (Shrestha and Orchard, 2018). In this framework, during forward propagation, the spike threshold uses a Heaviside function $H(x)$, while during back-propagation it employs the smooth surrogate
$\frac{\mathrm{d}H(x)}{\mathrm{d}x}\mathrel{\stackrel{{\scriptstyle\triangle}}{{=}}}\frac{1}{(\alpha\cdot|x|+1)^{2}},\quad\alpha=100,$ (1)
following the general form proposed by Neftci et al. (2019). Unless otherwise noted, this surrogate gradient is used consistently across all models.
During training, classification is supervised with a ProbSpikes cross-entropy loss on the output membrane potentials, and at inference, the predicted label is given by the index of the output neuron that fired the most spikes over the simulation time. ProbSpikes loss computes cross-entropy on normalized spike counts rather than membrane potentials: the spikes from each output neuron over the time window are converted into a probability distribution and compared with the one-hot target. This encourages the correct class to emit the largest fraction of spikes (Shrestha et al., 2022).
@@ 2.3 Delay Module Enhances Cross-Channel ISI Code Learning
This difference between the models with and without delays at longer intervals may be a result of the fact that the membrane time constant is more constrained in having to be long enough to integrate information across all timescales that are present, but short enough to be sensitive to differences on those timescales. By contrast, delays are not so constrained, and are therefore easier to adapt to tasks in which a wider range of timescales are present. We will return to this in section˜3.
@@ 2.4 Surrogate GD Enables Coincidence-Based Temporal Code Learning
In this task, each sample consists of $N=60$ input neurons evenly divided into three groups (2(b)). Time is divided into windows of length $w=200$ time steps, and in each window a group is either ON (high mean firing rate $\mu_{\text{on}}$) or OFF (low mean firing rate $\mu_{\text{off}}$). The class identity is defined by which two of the three groups fire with synchronous patterns (in every window, both are ON or both OFF) while the third inverts (OFF when the other two groups are ON, and vice versa). For example, Class A corresponds to synchrony between Groups 1 and 2, with Group 3 out of synchrony. To eliminate rate-based cues, ON/OFF group activity toggles randomly in each window, and so the total spike count is balanced across groups and samples, meaning that firing rate gives no information about class. Spike counts are Poisson distributed, and spikes are randomly allocated to times in the central 50% of each window. To control task difficulty, we interpolate between ON/OFF spike count distributions via a synchrony overlap factor $\lambda\in[0,1]$:
@@ 3 Validation of Spike Timing Learning on Speech Datasets
We have seen that surrogate gradient descent can learn to extract various forms of temporal structure (synchrony and ordered/unordered time differences within/across channels) in controlled conditions. We now test whether surrogate gradient–trained SNNs can exploit spike timing information in more realistic sensory tasks. To this end, we evaluate them on SHD and SSC, which consist of samples of single spoken words from 20/35 classes (respectively) converted to spike trains via a simulated cochlea model (Cramer et al., 2020). These datesets exhibit complex spatiotemporal structure but also strong spike-count variation that allows rate-based models to perform surprisingly well. To disentangle rate and timing contributions, we construct two timing-normalized dataset variants (fig.˜3) that progressively remove rate information while preserving each sample’s temporal structure.
@@ 3.1 Models
All experiments in this section use a two-layer feedforward SNN (fig.˜4) implemented in the SLAYER framework (Shrestha and Orchard, 2018), with 128 neurons per layer and the Spikemax loss function (Shrestha et al., 2022). We compare two variants: a baseline model trained by surrogate gradient descent (SGD), which learns synaptic weights and an enhanced version (SGD-delay) that additionally learns axonal delays (Sun et al., 2023). Predictions are obtained from the neuron with the highest spike count over the simulation window.
We also compare to the performance of a multilayer perceptron (MLP) using a vector of spike counts for each neuron as input. This serves as a proxy for the performance level attainable using only firing rates, and not spike times. The MLP comprises two fully connected layers with 256 hidden units and ReLU activation, mapping the spike count vector to a linear readout over all output classes.
@@ 3.2 Construction of Timing-Normalized SHD and SSC Datasets
While SHD and SSC exhibit complex spatiotemporal spike patterns (Cramer et al., 2020), spike count vectors vary significantly between classes and consequently, even simple multilayer perceptrons trained only on per‑sample spike counts can achieve nontrivial accuracy (for example, around 50% for SHD; fig.˜5). This makes it hard to determine whether SNNs truly leverage precise timing or simply exploit rate‑based cues.
@@
To address this, we construct two timing-normalized variants of the SHD and SSC datasets, each with an increasing amount of rate-based information stripped away. These variants aim to eliminate spike count confounds while preserving discriminative temporal structure. The normalization process follows a two-stage strategy:
@@ 3.3 Surrogate GD Enables SNNs to Exploit Spike Timing in Realistic Datasets
Figure 4: Comparison of SNN architectures trained with surrogate gradient descent (SGD) and SGD-delay (with delays). The top row shows a model with no delay components. The bottom row inserts learnable axonal delays between layers.
We evaluate SGD and SGD-delay on each variant (whole/part/norm), on both the SHD and SSC datasets, across the full range of spike timing perturbations $f\in[0,1]$ (fig.˜5). In every case where spike timing information has not been fully removed by perturbation ($f<1$), the spiking neural network model outperforms the MLP trained on the spike counts (which in turn give chance level performance for the Norm variant of each dataset). As $f\rightarrow 1$ the performance of the SNNs approaches the performance of the MLP using firing rates only. This confirms that surrogate gradient based-learning is not limited to learn rate-based features but can indeed learn from precise spike timing, both with and without explicit delay modeling.
@@ 4.3 Delay-Based Networks Use More Cross-Channel Timing Information
4.4 Delay-Based Networks Behave More Like Humans Under Time Reversal
Humans are highly sensitive to the causal structure of sensory input—e.g., speech played backward becomes unintelligible (Saberi and Perrott, 1999). To examine whether SNNs exhibit similar temporal sensitivity, we conduct a time reversal experiment by reversing spike trains within their active time windows (fig.˜7(a)).
@@ 4.4 Delay-Based Networks Behave More Like Humans Under Time Reversal
We train on the original data sets and test on the time-reversed versions. This makes sense when comparing to humans, since we do not listen to a lot of time reversed speech (even if we are big Twin Peaks fans). It also leaves all single neuron and synchrony timing cues unchanged, only affecting temporal order / causal cues, and therefore allows us to assess the extent to which the models make use of these different types of temporal cues. We also test at a variety of levels of perturbation $f$ (as in sections˜2 and 3).

### v3

@@ Abstract
  • The surrogate gradient descent algorithm enabled spiking neural networks to be trained to carry out challenging sensory processing tasks, an important step in understanding how spikes contribute to neural computations. However, it is unclear the extent to which these algorithms fully explore the space of possible spiking solutions to problems. We investigated whether spiking networks trained with surrogate gradient descent can learn to make use of information that is only encoded in the timing and not the rate of spikes. We constructed synthetic datasets with a range of types of spike timing information (interspike intervals, spatio-temporal spike patterns or polychrony, and coincidence codes). We find that surrogate gradient descent training can extract all of these types of information. In more realistic speech-based datasets, both timing and rate information is present. We therefore constructed variants of these datasets in which all rate information is removed, and find that surrogate gradient descent can still perform well. We tested all networks both with and without trainable axonal delays. We find that delays can give a significant increase in performance, particularly for more challenging tasks. To determine what types of spike timing information are being used by the networks trained on the speech-based tasks, we test these networks on time-reversed spikes which perturb spatio-temporal spike patterns but leave interspike intervals and coincidence information unchanged. We find that when axonal delays are not used, networks perform well under time reversal, whereas networks trained with delays perform poorly. This suggests that spiking neural networks with delays are better able to exploit temporal structure. To facilitate further studies of temporal coding, we have released our modified speech-based datasets.
@@ 1 Introduction
Spiking Neural Networks (SNNs) are biologically inspired models that compute via discrete, sparse spikes, rather than continuous activations (maass1997networks). This event‑driven framework not only captures rich temporal patterns (such as inter‑spike intervals and cross‑neuron synchrony) but also powers energy‑efficient neuromorphic hardware (painkras2013spinnaker; akopyan2015truenorth; davies2018loihi; orchard2021efficient; su2024ultra).
@@
Despite their promise, a persistent challenge remains: how to convert precise spike timing into effective learning signals. Surrogate Gradient Descent (Surrogate GD) provides a practical solution by replacing the non‑differentiable spike with a smooth surrogate, thereby allowing backpropagation through time (neftci2019surrogate). Subsequent work has improved its efficiency and adaptiveness by enforcing sparser gradients (perez2021sparse) and by developing more adaptive surrogate functions (wang2023adaptive). Surrogate GD–trained networks have matched or outperformed rate‑based recurrent neural networks on several neuromorphic benchmarks (shrestha2018slayer; 10757311; dampfhoffer2024neuromorphic; hammouamri2024learning; sun2025towards), but it remains unclear whether these successes stem from genuine temporal coding or from reliance on coarse firing‑rate features.
@@
To address this ambiguity, researchers have injected various temporal inductive biases: heterogeneous time constants (fang2021incorporating; perez2021neural; habashy2024adapting; yu2025neuronal); learnable axonal delays (shrestha2022spikemax; sun2022axonal; 10181778; sun2025algorithm); synaptic delays (zhang2020supervised; yu2022improving; sun2024delay; hammouamri2024learning; queant2025delrec); and dendritic delay models (d2024denram; zheng2024temporal). Yet even with these enhancements, there has been no systematic test of whether or not Surrogate GD finds networks that make use of precise timing.
@@
This uncertainty leads to a central question: can Surrogate GD-trained SNNs learn and depend on spike timing codes beyond rate statistics? This question is important because if they were just using rate-based information, then (a) they are not making use of all the information that is available, and are therefore less efficient than they could be, (b) it limits their effectiveness as a model of spike-based computation in the brain, and (c) it raises the question of whether or not there is an advantage of using such expensive training methods compared to simply approximating rate-based networks. Previous work has demonstrated that surrogate gradient methods can be trained to reproduce prescribed spiking patterns or low-dimensional spike manifolds (zenke2018superspike). However, such approaches typically impose strong constraints on the input or target spike structure (zenke2021remarkable). To address this central question, we used both synthetic datasets as well as more complicated real-world benchmarks. First, we construct synthetic benchmark tasks that isolate the contribution of inter-spike interval (ISI), coincidence, and cross-channel inter-spike interval (CCISI, a type of spatio-temporal spike pattern or polychrony) under controlled spike counts, building upon prior work that highlights the ability of SNNs to leverage temporal coding for classification (bohte2002error; mirsadeghi2021stidi). Then, we evaluate models on datasets with real-world temporal complexity (cramer2020heidelberg), introducing spike count-normalized variants to remove rate confounds. We assess robustness under biologically inspired perturbations: Gaussian jitter applied per-spike and per-neuron (amarasingham2012conditional; agmon2012novel), and spike deletions (mainen1995reliability; chen2022reliability). Finally, we analyse the effect of time reversal that selectively disrupts spike patterns while leaving coincidences and ISIs unchanged.
@@
Our findings show that Surrogate GD-trained SNNs can extract fine-grained temporal structure even without spike rate information, including intra-neuron ISIs, cross-neuron synchrony or coincidence, and cross-channel causal ISI. Delay learning further enhances the model’s ability to exploit such timing features, especially for long-range causal gaps that exceed the membrane integration window controlled by the time constant. Across perturbations we observe distinct degradation patterns that reveal both vulnerabilities and robustness in the model’s temporal encoding. Notably, under time reversal, Surrogate GD-trained SNNs not only leverage reversal-invariant features like ISI and coincidence, but also capture causal dependencies in spike timing (as in CCISI), indicating a layered temporal representation. The implementation details are available at https://github.com/neural-reckoning/temporal-shd.
@@ 2.1 Models
Training is implemented in the SLAYER framework (shrestha2018slayer). In this framework, during forward propagation, the spike threshold uses a Heaviside function $H(x)$, while during back-propagation it employs the smooth surrogate
$\frac{\mathrm{d}H(x)}{\mathrm{d}x}\triangleq\frac{1}{(\alpha\cdot|x|+1)^{2}},\quad\alpha=100,$ (1)
following the general form proposed by (neftci2019surrogate). Unless otherwise noted, this surrogate gradient is used consistently across all models. This surrogate derivative provides a simple mechanism by which gradients become sensitive to spike timing. When an input spike arrives slightly earlier or later, the postsynaptic membrane potential crosses threshold at a different time, and the surrogate derivative in Eq. (1) produces a correspondingly shifted gradient. In combination with learned delays, this creates an effective temporal receptive field: the network strengthens weights or adjusts delays for spike timings that consistently support correct classification.
During training, classification is supervised with a Spikemax loss on the output spikes over the simulation time (shrestha2022spikemax). In the synthetic CCISI and ISI experiments, both the membrane time constant $\tau$ and the axonal delays are treated as fully learnable parameters and optimized jointly with all network weights using the surrogate gradient method (neftci2019surrogate; sun2022axonal).
@@ Results.
This difference between the models with and without delays at longer intervals may be a result of the fact that the membrane time constant is more constrained in having to be long enough to integrate information across all timescales that are present, but short enough to be sensitive to differences on those timescales. By contrast, delays are not so constrained, and are therefore easier to adapt to tasks in which a wider range of timescales are present. We will return to this in section˜3. This distinction between integration timescales and explicit temporal alignment is consistent with prior work showing that learnable axonal delays provide a complementary mechanism to heterogeneous membrane time constants for temporal coding (sun2023learnable; sun2025exploiting).
@@ 2.4 Surrogate GD Enables Coincidence-Based Temporal Code Learning
In this task, each sample consists of $N=60$ input neurons evenly divided into three groups (fig.˜2(b)). Time is divided into windows of length $w=200$ time steps, and in each window a group is either ON (high mean firing rate $\mu_{\text{on}}$) or OFF (low mean firing rate $\mu_{\text{off}}$). The class identity is defined by which two of the three groups fire with synchronous patterns (in every window, both are ON or both OFF) while the third inverts (OFF when the other two groups are ON, and vice versa). For example, Class A corresponds to synchrony between Groups 1 and 2, with Group 3 out of synchrony. To eliminate rate-based cues, ON/OFF group activity toggles randomly in each window, and so the total spike count is balanced across groups and samples, meaning that firing rate gives no information about class. Spike counts are Poisson distributed, and spikes are randomly allocated to times in the central 50% of each window. To control task difficulty, we interpolate between ON/OFF spike count distributions via a synchrony overlap factor $\lambda\in[0,1]$:
@@ 3 Validation of Spike Timing Learning on Speech Datasets
We have seen that surrogate gradient descent can learn to extract various forms of temporal structure (synchrony and ordered/unordered time differences within/across channels) in controlled conditions. We now test whether surrogate gradient–trained SNNs can exploit spike timing information in more realistic sensory tasks. To this end, we evaluate them on SHD and SSC, which consist of samples of single spoken words from 20/35 classes (respectively) converted to spike trains via a simulated cochlea model (cramer2020heidelberg). Following the standard SHD preprocessing, each sample is discretized into 100 temporal bins sun2023learnable.These datesets exhibit complex spatiotemporal structure but also strong spike-count variation that allows rate-based models to perform surprisingly well. To disentangle rate and timing contributions, we construct two timing-normalized dataset variants (fig.˜3) that progressively remove rate information while preserving each sample’s temporal structure.
@@ 3.1 Models
All experiments in this section use a two-layer feedforward SNN (fig.˜4) implemented in the SLAYER framework (shrestha2018slayer), with 128 neurons per layer and the Spikemax loss function (shrestha2022spikemax). We compare two variants: a baseline model trained by surrogate gradient descent (SGD), which learns synaptic weights and an enhanced version (SGD-delay) that additionally learns axonal delays (sun2023learnable). Predictions are obtained from the neuron with the highest spike count over the simulation window.
We also compare to the performance of a multilayer perceptron (MLP) using a vector of spike counts for each neuron as input. This serves as a proxy for the performance level attainable using only firing rates, and not spike times. The MLP comprises two fully connected layers with 128 hidden units and ReLU activation, mapping the spike count vector to a linear readout over all output classes.
For the SHD and SSC datasets, we use 200 simulation time steps. Axonal delays are initialised uniformly in the range $[0,1]$, and an adaptive delay threshold is applied to regulate them during training 10094768. All temporal perturbations (jitter, deletion, reversal) and class-balancing procedures for the Part and Norm variants follow a fixed protocol.
@@ 3.2 Construction of Timing-Normalized SHD and SSC Datasets
While SHD and SSC exhibit complex spatiotemporal spike patterns (cramer2020heidelberg), spike count vectors vary significantly between classes and consequently, even simple multilayer perceptrons trained only on per‑sample spike counts can achieve nontrivial accuracy (for example, around 50% for SHD; fig.˜5). This makes it hard to determine whether SNNs truly leverage precise timing or simply exploit rate‑based cues.
@@
To address this, we construct two timing-normalized variants of the SHD and SSC datasets, each with an increasing amount of rate-based information stripped away. These variants aim to eliminate spike count confounds while preserving discriminative temporal structure. The normalization process follows a two-stage strategy. For clarity, we briefly note that the following subsections describe this process in two stages: Section 3.2 first introduces the neuron and sample filtering procedure used to obtain the Part variant, and Section 3.2.2 then explains how the Norm variant is produced by subsampling spike trains to remove all rate-based information while preserving timing.
@@ 3.3 Surrogate GD Enables SNNs to Exploit Spike Timing in Realistic Datasets
Figure 4: Comparison of Feedforward SNN architectures trained without delays (SGD, top) and with delays (SGD-delay, bottom). Both models adapt the surrogate gradient descent methods. The top row shows a model with no delay components. The bottom row inserts learnable axonal delays between layers. Dashed outlines are used purely as a visual aid to highlight the architectural comparison between the two variants.
We evaluate SGD and SGD-delay on each variant (whole/part/norm), on both the SHD and SSC datasets, across the full range of spike timing perturbations $f\in[0,1]$ (fig.˜5). Here, $f$ is the same perturbation factor used in the synthetic tasks. All reported accuracies are averaged over 5 independent runs with different random seeds; shaded regions in the figures indicate mean $\pm$ standard deviation across runs. In every case where spike timing information has not been fully removed by perturbation ($f<1$), the spiking neural network model outperforms the MLP trained on the spike counts (which in turn give chance level performance for the Norm variant of each dataset). As $f\rightarrow 1$ the performance of the SNNs approaches the performance of the MLP using firing rates only. This confirms that surrogate gradient based-learning is not limited to learn rate-based features but can indeed learn from precise spike timing, both with and without explicit delay modeling.
@@ 4.3 Delay-Based Networks Use More Cross-Channel Timing Information
4.4 Delay-Based Networks Are More Sensitive To Time Reversal
Humans are highly sensitive to the causal structure of sensory input -— e.g., speech played backward becomes unintelligible (saberi1999cognitive). To examine whether SNNs exhibit similar temporal sensitivity, we conduct a time reversal experiment by reversing spike trains within their active time windows (fig.˜7(a)). Note that in our simple model here, time reversal refers strictly to reversing the temporal order of spikes within each spike train, while preserving spike counts, neuron identities, and the underlying time discretization. A more realistic model of humans listening to time-reversed speed would be to produce new spike trains by running the time-reversed speech through the auditory model, but this would not allow for us to control for keeping the spike counts and ISI distributions identical.
@@ 4.4 Delay-Based Networks Are More Sensitive To Time Reversal
We train on the original data sets and test on the time-reversed versions. This makes sense when comparing to humans, since we do not listen to a lot of time reversed speech (even if we are big Twin Peaks fans). It also leaves all single neuron and synchrony timing cues unchanged, only affecting temporal order / causal cues (including CCISI), and therefore allows us to assess the extent to which the models make use of these different types of temporal cues. We also test at a variety of levels of perturbation $f$ (as in sections˜2 and 3).